What does the fact that this particular graph is linear indicate about the acceleration? What is the acceleration of Object 3?

Exercise 2 Report - Word References Mailings Review View Easy Document Creator A.y.A- 를 15-12-B- 11 Heading 111Normal! Subtitle Paragraph Position versus Time 2 (x vs tA2) for Object 3 45 40 35 30 25 20 15 10 40.3 : 34.9 y 0.0628x + 03256 25.5 20.1 15.2 118 8.4 4.3 5 061 100 200300 400 500 600 tA2 (sec #2)
Answer #1

straight line indicates that acceleration of the object is constant.

x = 1/2*a*t2

so acceleration = 2*slope of line = 2*0.0628 = 0.1256 m/s2